OpenStudy (anonymous):

Solve each of the following equations. Give exact answers. a) 4x-3 = 7 b) 15x2+x-6 = 0 c) 3x2-8x-1 = 0

OpenStudy (anonymous):

Do you have anymore information?

OpenStudy (anonymous):

No that's it.. but its a topic on logarithm and exponential equation

sammixboo (sammixboo):

Ok, let's do A. 4x - 3 = 7 4/4x - 3 = 7/4 First, tell me what is 7 divided by 4 (7/4)

OpenStudy (anonymous):

7/4 = 1.75

OpenStudy (anonymous):

BroTip: It's always better to leave it in fractional form

OpenStudy (anonymous):

Yeah.. rather than solving it outright. It makes it a little easier my friend said to me as well

OpenStudy (anonymous):

do you still need help ?

OpenStudy (anonymous):

if not you should close this post

OpenStudy (anonymous):

quadratic formula. ax^2 + bx + c = 0. x = [-b + sqrt(b^2 - 4ac)] / (2a) or x = [-b - sqrt(b^2 - 4ac)] / (2a).

OpenStudy (anonymous):

where a, b and c are some constants (in short, some numbers which will not change), you will have.

OpenStudy (anonymous):

3x^2 - 8x - 1 = 0 3(x^2 - (8/3)x) - 1 = 0 3(x - (8/6))^2 - 3(64/36) - 1 = 0 3(x - (8/6))^2 - 16/3 - 1 = 0 3(x - (8/6))^2 - 19/3 = 0 3(x - (8/6))^2 = 19/3 (x - (8/6))^2 = 19/9 x - (8/6) = sqrt(19/9) or = -sqrt(19/9) x = sqrt(19/9) + (8/6) or x = -sqrt(19/9) + (8/6) x = sqrt(19/9) + (4/3) or x = -sqrt(19/9) + (4/3)

OpenStudy (anonymous):

3x^2 - 8x - 1 = 0 3x^2 + (-8x) + (-1) = 0 a = 3, b = -8 and c = -1

OpenStudy (anonymous):

using quadratic formula. x = [8 + sqrt((-8)^2 - 4(3)(-1))] / (2)(3) or x = [8 - sqrt((-8)^2 - 4(3)(-1))] / (2)(3) x = [8 + sqrt(64 + 12)] / 6 or x = [8 - sqrt(64 + 12)] / 6 x = (8/6) + sqrt(76)/6 or x = (8/6) - sqrt(76/6) x = (4/3) + [sqrt(4)sqrt(19)] / [(2)(3)] or x = (4/3) - [sqrt(4)sqrt(19)] / [(2)(3)] x = (4/3) + [2sqrt(19)] / [(2)(3)] or x = (4/3) - [2sqrt(19)] / [(2)(3)] x = (4/3) + sqrt(19)/3 or x = (4/3) - sqrt(19)/3 x = (4/3) + sqrt(19)/sqrt(9) or x = (4/3) - sqrt(19)/sqrt(9) x = (4/3) + sqrt(19/9) or x = (4/3) - sqrt(19/9).